CPS 130 Homework 20 Minimum Spanning Tree

due Fri June 21st

Write and justify your answers in the space provided.^{1}

1. [CLRS 23.2-4]

2. [CLRS 23-4]

1Collaboration is allowed, even encouraged, provided that the names of the collaborators are listed along with the solutions. Students must write up the solutions on their own.

2